# Non-equilibrium dynamics and atom-pair coherence in strongly interacting Bose-Fermi mixtures

### v1

@@ III.2 Quasi-stationary approximation
Now that we have established the pairing physics within the mixture, we can examine whether this physics is correctly contained within the equations of motion formulated in Sec. II. For that we will focus on the dynamics of the cumulants $L_{\mathbf{Q},\frac{m_{F}}{M}\mathbf{Q}-\mathbf{q}}^{*}$ and $\chi_{\mathbf{Q}}$, interpreted as open and closed-channel components of the embedded two-body wave function. This is most easily shown by taking Eqs. (II.2) and (II.2) and neglecting inhomogeneous source terms from the BEC and bosonic pairing matrix. Then we adopt the quasi-stationary ansatz,
@@
where $\Delta E_{\mathbf{Q}}=E_{\mathbf{Q}}-\varepsilon_{\mathbf{Q}}^{F}$ quantifies the energy difference between both sides of the fundamental scattering diagram in Fig. 1(a), renormalized such that the energy of the molecule equals the dressed dimer energy $E_{\mathbf{Q}}$. Eqs. (II.2) and (II.2) can now be rewritten as eigenvalue equations,
@@
for the pair energy $E_{\mathbf{Q}}$. Eqs. (III.2) and (25) are equivalent to the momentum-space two-body Schrödinger equation for the open- and closed-channel wave functions $\Psi_{\mathbf{Q},\mathbf{q}}$ and $\Phi_{\mathbf{Q}}$ respectively, with the two-body interaction enhanced by the factor $(1+n_{\frac{m_{B}}{M}\mathbf{Q}+\mathbf{q}}^{B}-n_{\frac{m_{F}}{M}\mathbf{Q}-\mathbf{q}}^{F})$ (see Fig. 1b). Assuming no bosonic excitations $n_{\mathbf{k}}^{B}=0$ and an ideal Fermi sea $n_{\mathbf{k}}^{F}=\Theta(k_{F}-k)$, the solutions $E_{\mathbf{Q}}$ of Eqs. (III.2) and (25) are given by Eq. (III.1), provided that we take the limit $n_{0}^{B}\rightarrow 0$. This proves the correct embedding of the two-body Schrödinger equation for interactions between fermions and noncondensed bosons. The appropriate inclusion of the mean field effects due to the BEC originate from the source terms scaling with $\psi$, including an additional mutual coupling between $L_{\mathbf{Q},\frac{m_{F}}{M}\mathbf{Q}-\mathbf{q}}^{*}$ and $M_{\mathbf{Q},\frac{m_{F}}{M}\mathbf{Q}-\mathbf{q}}^{*}$. Due to the nonlinear nature of the associated differential equations however, these effects can only be confirmed numerically.
@@ III.3 Atom-pair coherence
We observe that, for the full range of scattering lengths we study, the oscillations in the population of the condensate match well with the energies of the embedded pair, supporting our hypothesis that the oscillations seen in Fig. 4(a) originate from a coherent exchange of bosons between the BEC and the paired state. Especially interesting is that this coherence remains robust even for negative scattering lengths, where the exchange occurs between the BEC and the anomalous dimer in the Fermi sphere. This is a unique feature of the Bose-Fermi mixture, as a similar coherence in identical bose gases is not possible due to the absence of the Feshbach dimer for attractive interactions. At the crossing point $(k_{F}a_{BF})^{-1}=2/\pi$, where the absolute energies of the Feshbach and anomalous dimer states are equal, we observe that the competing exchange between the two states leads to a slight increase of the oscillation frequency, suggesting some interference effects. For small negative scattering lengths, we find that the oscillations become increasingly incoherent, in accordance with the merging of the anomalous dimer with the two-body scattering continuum, to which the coherent oscillations show rapid dephasing as a function of time.
@@ IV.1 Condensate depletion and molecule production
First, we examine in Figs. 7 and 8 the depletion of the BEC and associated generation of bosonic excitations and fermionic molecules, which we find depends strongly on both the width of the Feshbach resonance and the mass ratio. We can quantify this observation more precisely by solving the cumulant equation of motion for $\chi_{\mathbf{k}}$ in the early time limit, capturing the scattering diagram in Fig. 1(a). Equation the populations in Eq. (II.2) with their initial ($t=0$) values, $\psi=\sqrt{n_{B}}$, $n_{\mathbf{k}}^{F}=\Theta(k_{F}-k)$ and $n_{\mathbf{k}}^{M}=0$, the resulting expression can be integrated exactly. Upon substituting the result into Eqs. (9), (II.2) and (II.2), one finds
@@ IV.2 Momentum distribution
At longer times following the quench, the dynamics set by the diagram in Fig. 1(a) compete with the higher order scattering diagram in Fig. 1(b), which generates bosonic excitations at $k>0$ and fermionic excitations at $k>k_{F}$, see Figs. 9(a,c). Contrary to the unitary Bose gas, where $n_{\mathbf{k}}^{B}$ is typically a monotonously decreasing function of momentum Colussi et al. (2020), the generation of bosonic excitations at momenta $k\ll k_{F}$ is suppressed by Pauli blockade from the Fermi sphere, which ensures that such low-momentum excitations can only be generated by high-momentum molecules with $Q\approx k_{F}$. In the bosonic distribution function, this effect competes with the universal decay of $n_{\mathbf{k}}^{B}$ at large momentum $k\gg k_{F}$ as set by the two-body contact Werner and Castin (2012), thus inducing a peak in $n_{\mathbf{k}}^{B}$ at $k\approx k_{F}$ where excitations can originate from all molecules inside the molecular Fermi sea. Hence, the same physics that leads to the appearance of an anomalous dimer state in the two-body spectrum discussed in Sec. III, also imprints the presence of the Fermi sea on the bosonic excitation density following a quench, as can be clearly seen in Fig. 9(a).
@@ V CONCLUSION
Looking ahead, our exploratory work suggests several interesting directions for future research, chief amongst these being the important question whether the deep quench dynamics that we have examined can also be seen in an experiment, where non-coherent effects such as particle loss from the trap are also likely to affect observations. Furthermore, it is known from the unitary Bose gas that three-body correlations, enhanced by the Efimov effect, can strongly affect the heating dynamics of the gas for $t>t_{F}$ Musolino et al. (2022); van de Kraats et al. (2024). As the Efimov effect also manifests in Bose-Fermi mixtures Sun and Cui (2019), it will be interesting to examine such correlations also in this system. Similarly, open questions remain regarding the effects of mediated Bose-Bose interactions in a dynamical setting DeSalvo et al. (2019); Edri et al. (2020); Shen et al. (2024). Finally, our model may provide a new tool for studying the out-of-equilibrium dynamics of impurities, by taking the limit in which either the bosonic or fermionic density is very small, thus recovering the canonical problem of the Fermi or Bose polaron, which has recently seen significant experimental interest Fritsche et al. (2021); Skou et al. (2021, 2022); Baroni et al. (2024).

### v2

@@ III.2 Quasi-stationary approximation
Now that we have established the pairing physics within the mixture, we can examine whether this physics is correctly contained within the equations of motion formulated in Sec. II. For that we will focus on the dynamics of the cumulants $L_{\mathbf{Q},\frac{m_{F}}{M}\mathbf{Q}-\mathbf{q}}^{*}$ and $\chi_{\mathbf{Q}}$, interpreted as open and closed-channel components of the embedded two-body wave function. This is most easily shown by taking Eqs. (II.2) and (II.2) and neglecting inhomogeneous source terms from the BEC and bosonic pairing matrix by setting $\psi=\kappa_{\mathbf{k}}=0$ Kira (2015). We then obtain the coupled equations of motion,
$i\dot{\chi}_{\mathbf{Q}}$ $=$ $(\varepsilon_{\mathbf{Q}}^{M}+\nu-\varepsilon_{\mathbf{Q}}^{F})\chi_{\mathbf{Q}}$
$+$ $\frac{g}{\sqrt{V}}\sum_{\mathbf{q}}\xi_{\mathbf{q}}L_{\mathbf{Q},\frac{m_{F}}{M}\mathbf{Q}-\mathbf{q}}^{*},$
$i\dot{L}_{\mathbf{Q},\frac{m_{F}}{M}\mathbf{Q}-\mathbf{q}}^{*}$ $=$ $(\varepsilon_{\mathbf{Q}}^{M}+\varepsilon_{\mathbf{q}}^{R}-\varepsilon_{\mathbf{Q}}^{F})L_{\mathbf{Q},\frac{m_{F}}{M}\mathbf{Q}-\mathbf{q}}^{*}$
$+$ $\frac{g}{\sqrt{V}}\xi_{\mathbf{q}}(1+n_{\frac{m_{B}}{M}\mathbf{Q}+\mathbf{q}}^{B}-n_{\frac{m_{F}}{M}\mathbf{Q}-\mathbf{q}}^{F})\chi_{\mathbf{Q}}.$
Here we have assumed that the dynamics of the densities are much slower than the dynamics of $L_{\mathbf{Q},\frac{m_{F}}{M}\mathbf{Q}-\mathbf{q}}^{*}$ and $\chi_{\mathbf{Q}}$, such that they may be approximated as constant, also known as the quasi-stationary approximation Colussi et al. (2018). To solve Eqs. (III.2) we adopt the ansatz,
@@
where $\Delta E_{\mathbf{Q}}=E_{\mathbf{Q}}-\varepsilon_{\mathbf{Q}}^{F}$ quantifies the energy difference between both sides of the fundamental scattering diagram in Fig. 1(a), renormalized such that the energy of the molecule equals the dressed dimer energy $E_{\mathbf{Q}}$. Eqs. (III.2) can now be rewritten as eigenvalue equations,
@@
for the pair energy $E_{\mathbf{Q}}$. Eqs. (III.2) and (27) are equivalent to the momentum-space two-body Schrödinger equation for the open- and closed-channel wave functions $\Psi_{\mathbf{Q},\mathbf{q}}$ and $\Phi_{\mathbf{Q}}$ respectively, with the two-body interaction again enhanced by the factor $(1+n_{\frac{m_{B}}{M}\mathbf{Q}+\mathbf{q}}^{B}-n_{\frac{m_{F}}{M}\mathbf{Q}-\mathbf{q}}^{F})$. Note here that the single-particle momenta match the diagram in Fig. 1b.
Eqs. (III.2) and (27) constitute a set of linear algebraic equations, which may be directly diagonalized to obtain the pair energies $E_{\mathbf{Q}}$. Assuming no bosonic excitations $n_{\mathbf{k}}^{B}=0$ and an ideal Fermi sea $n_{\mathbf{k}}^{F}=\Theta(k_{F}-k)$, one then finds that $E_{\mathbf{Q}}$ obeys Eq. (III.1), provided that we also take the limit $n_{0}^{B}\rightarrow 0$. This proves the correct embedding of the two-body Schrödinger equation for interactions between fermions and noncondensed bosons. The appropriate inclusion of the mean field effects due to the BEC originate from the source terms scaling with $\psi$, including an additional mutual coupling between $L_{\mathbf{Q},\frac{m_{F}}{M}\mathbf{Q}-\mathbf{q}}^{*}$ and $M_{\mathbf{Q},\frac{m_{F}}{M}\mathbf{Q}-\mathbf{q}}^{*}$. Due to the nonlinear nature of the associated differential equations however, these effects can only be confirmed numerically, as we will do in the next section.
@@ III.3 Atom-pair coherence
We observe that, for the full range of scattering lengths we study, the oscillations in the population of the condensate match well with the energies of the embedded pair, supporting our hypothesis that the oscillations seen in Fig. 4(a) originate from a coherent exchange of bosons between the BEC and the paired state. Especially interesting is that this coherence remains robust even for negative scattering lengths, where the exchange occurs between the BEC and the anomalous dimer in the Fermi sphere. This is a unique feature of the Bose-Fermi mixture, as a similar coherence in identical bose gases is not possible due to the absence of the Feshbach dimer for attractive interactions. At the crossing point $(k_{F}a_{BF})^{-1}=2/\pi$, where the absolute energies of the Feshbach and anomalous dimer states are equal, we observe that the competing exchange between the two states leads to enhanced interference, resulting in a slight increase of the dominant oscillation frequency. For small negative scattering lengths, we find that the oscillations become increasingly incoherent, in accordance with the merging of the anomalous dimer with the two-body scattering continuum, to which the coherent oscillations show rapid dephasing as a function of time.
@@ IV.1 Condensate depletion and molecule production
First, we examine in Figs. 7 and 8 the depletion of the BEC and associated generation of bosonic excitations and fermionic molecules, which we find depends strongly on both the width of the Feshbach resonance and the mass ratio. We can quantify this observation more precisely by solving the cumulant equation of motion for $\chi_{\mathbf{k}}$ in the early time limit, capturing the scattering diagram in Fig. 1(a). Equating the populations in Eq. (II.2) with their initial ($t=0$) values, $\psi=\sqrt{n_{B}}$, $n_{\mathbf{k}}^{F}=\Theta(k_{F}-k)$ and $n_{\mathbf{k}}^{M}=0$, the resulting expression can be integrated exactly. Upon substituting the result into Eqs. (9), (II.2) and (II.2), one finds
@@ IV.2 Momentum distribution
At longer times following the quench, the dynamics set by the diagram in Fig. 1(a) compete with the higher order scattering diagram in Fig. 1(b), which generates bosonic excitations at $k>0$ and fermionic excitations at $k>k_{F}$, see Figs. 9(a,c). Contrary to the unitary Bose gas, where $n_{\mathbf{k}}^{B}$ is typically a monotonously decreasing function Colussi et al. (2020), the generation of bosonic excitations at momenta $k\ll k_{F}$ is suppressed by Pauli blockade from the Fermi sphere, which ensures that such low-momentum excitations can only be generated by high-momentum molecules with $Q\approx k_{F}$. In the bosonic distribution function, this effect competes with the universal decay of $n_{\mathbf{k}}^{B}$ at large momentum $k\gg k_{F}$ as set by the two-body contact Werner and Castin (2012), thus inducing a peak in $n_{\mathbf{k}}^{B}$ at $k\approx k_{F}$ where excitations can originate from all molecules inside the molecular Fermi sea. Hence, the same physics that leads to the appearance of an anomalous dimer state in the two-body spectrum discussed in Sec. III, also imprints the presence of the Fermi sea on the bosonic excitation density following a quench, as can be clearly seen in Fig. 9(a).
@@ V CONCLUSION
Looking ahead, our exploratory work suggests several interesting directions for future research, chief amongst these being the important question whether the deep quench dynamics that we have examined can also be seen in an experiment, where inite-size effects and non-coherent physics such as particle loss from the trap are also likely to affect observations. Furthermore, it is known from the unitary Bose gas that three-body correlations, enhanced by the Efimov effect, can strongly affect the heating dynamics of the gas for $t>t_{F}$ Musolino et al. (2022); van de Kraats et al. (2024). As the Efimov effect also manifests in Bose-Fermi mixtures Sun and Cui (2019), it will be interesting to examine such correlations also in this system. Similarly, open questions remain regarding the effects of mediated Bose-Bose interactions in a dynamical setting DeSalvo et al. (2019); Edri et al. (2020); Shen et al. (2024). Finally, our model may provide a new tool for studying the out-of-equilibrium dynamics of impurities, by taking the limit in which either the bosonic or fermionic density is very small, thus recovering the canonical problem of the Fermi or Bose polaron, which has recently seen significant experimental interest Fritsche et al. (2021); Skou et al. (2021, 2022); Baroni et al. (2024).